OpenStudy (anonymous):

What is the value of the x variable in the solution to the following system of equations? 3x + 2y = 8 2x + 3y = 2

OpenStudy (anonymous):

somebody help me

OpenStudy (lgbasallote):

3x + 2y = 8 2x + 3y =2 9x + 6y = 24 -4x -6y = -4 5x = 20 x = 4

OpenStudy (agentnao):

And if x equals 4, then y is -2